# Distributionally Robust Policy Learning under Concept Drifts

### v1

@@ 6 Discussion
In this paper, we study the policy learning problem under concept drift, where we propose a doubly-robust policy value estimator that is consistent and asymptotically normal, and then develop a minimax optimal policy learning algorithm, whose regret is $O_{p}(\kappa(\Pi)n^{-{1}/{2}})$ with a matching lower bound. Numerical results show that our learning algorithm outperforms the benchmark algorithm under concept drift. Future research could extend our results on pure concept drift to a more general setting where the concept drift adopts the same form as ours, but there is in addition an identifiable covariate shift as in Jin et al., 2022b .
@@ Acknowledgments
The authors would like to thank Miao Lu and Wenhao Yang for pointing out an error in an earlier draft of this manuscript and their helpful comments. Z.R. is supported by NSF grant DMS-2413135.
Appendix A Experiment details
In the data generating process, we choose $\beta$’s and $\sigma$’s to be
@@ Appendix A Experiment details
The underlying policy $\pi_{0}$ chooses actions with context $x$ according to the following rules:
@@
We generate $\mathcal{D}_{\text{train}}$ according to the procedure described above as the training dataset. We also generate 10,000 samples as our testing dataset $\mathcal{D}_{\text{test}}=\{i\in[10,000]:(X_{i},Y_{i}(1),Y_{i}(2),Y_{i}(3))\}$, which we use to estimate the true policy value.
@@
Appendix B Deferred proofs of the main results
B.1 Proof of Lemma 2.3
@@ Appendix B Deferred proofs of the main results
where the function $f_{\text{KL}}(x)=x\log x$ represents the KL divergence function. In (B.1), the first constraint reflects that $L$ is an likelihood ratio, and the second constraint corresponds to the KL divergence bound.
For notational simplicity, let $\mathbb{E}_{x}$ be the shorthand of $\mathbb{E}_{P_{Y\,|\,X}}[\cdot\,|\,X=x]$. By Theorem 8.6.1 of Luenberger, (1997), the Slater’s condition is satisfied and strong duality holds:
@@
B.2 Proof of Theorem 3.5
For notational simplicity, we drop the dependence on $P$ in $\mathbb{E}_{P}$ when the context is clear. The proof of Theorem 3.5 makes use of the following lemma, which establishes some useful properties of the optimizer $\bm{\theta}^{*}_{\pi}$. The proof of Lemma B.1 can be found in Appendix C.1.
Lemma B.1.
@@
By the first-order condition established in part (1) of Lemma B.1, we have
@@
where the last step is due to Equation (13). By Assumption 3.4, $\|\widehat{\bm{\theta}}^{(k)}_{\pi}-\bm{\theta}^{*}_{\pi}\|_{L_{\infty}}=o_{P}(1)$. Therefore, for any $\beta\in(0,1)$, there exists $N\in\mathbb{N}_{+}$ such that for $n\geq N$, $\|\widehat{\theta}_{\pi}^{(k)}-\theta^{*}_{\pi}\|_{L_{\infty}}\leq\min(\xi,\xi_{1})$. On the event that $\|\widehat{\theta}_{\pi}^{(k)}(x)-\theta^{*}_{\pi}(x)\|_{L_{\infty}}\leq\min(\xi,\xi_{1})$ by part (2) of Lemma B.1 and Jensen’s inequality, we have
@@
where the last step is due to part (3) of Lemma B.1. Combining the above results, we have that
@@
Above, step (i) follows from Jensen’s inequality and step (ii) from part (3) of Lemma B.1. Combining everything, we have that term (IV) is of rate $o_{P}(n^{-1/2})$.
@@
For any $a\in[M]$, if we take $\pi(x)\equiv a$, then by (1) of Lemma B.1,
@@
By (2) of Lemma B.1, for any $a\in[M]$, there exists a constant $\xi_{a}>0$ such that for any $\|\theta-\bm{\theta}^{*}_{a}(x)\|_{2}\leq\xi_{a}$
@@
By (3) of Lemma B.1, for any $a\in[M]$, there exists a constant $\xi_{1,a}>0$ such that for any $\|\bm{\theta}-\bm{\theta}^{*}_{a}\|_{L_{\infty}}\leq\xi_{1,a}$,
@@ B.3.2 The effective term
Above, the first inequality is because of the definition of $\sup$ and the second is due to the triangle inequality; the last step is due to the boundedness of $h$. Taking the supremum over all $\pi\in\Pi$ in (B.3.2), we have that
@@
By the bounded difference inequality (Wainwright,, 2019, Corollary 2.21), for any $t>0$,
@@
B.3.3 Bounding the negligible term
@@
Applying Lemma B.2 with $c_{i}=c_{1}(X_{i})$, we have that
@@
where the last step follows from part (1) of Lemma B.1. By Assumption 3.4, for any $\beta\in(0,1)$, there exists $N_{3}\in\mathbb{N}_{+}$ such that when $n\geq N_{3}$,
@@
Then by Lemma B.2, we have
@@
B.4 Proof of Theorem 4.6
We first state somes results from Si et al., (2023) that will be used in the proof. For any $p,q\in[0,1]$, define
@@
Lemma B.3 (Adapted from Lemma A17 of Si et al., (2023)).
@@
Note that our definition of $g_{\delta}(q)$ is slightly different from that in Si et al., (2023), so we include the proof of Lemma B.3 in Appendix C.4 for completeness.
@@
where step (i) uses that $g$ is non-decreasing (c.f. Cauchois et al., (2024, Proposition 1)); in step (ii), $\xi\in((1-\Delta)/2,(1+\Delta)/2)$, and step (iii) follows from Lemma B.3.
@@
(B.4)$\geq\sqrt{\frac{d}{n}}\frac{d2^{d-1}}{15d2^{d+2}}=\frac{1}{120}\times\sqrt{\frac{d}{n\varepsilon}}.$
Appendix C Proof of technical lemmas
C.1 Proof of Lemma B.1

### v2

@@ 6 Discussion
In this paper, we study the policy learning problem under concept drift, where we propose a doubly-robust policy value estimator that is consistent and asymptotically normal, and then develop a minimax optimal policy learning algorithm, whose regret is $O_{p}(\kappa(\Pi)n^{-{1}/{2}})$ with a matching lower bound. Numerical results show that our learning algorithm outperforms the benchmark algorithm under concept drift. We also note that our results on pure concept drift could be extended to a more general setting where the concept drift adopts the same form as ours, but there is in addition an identifiable covariate shift as in Jin et al. (2022b). The details can be found in Appendix E.
@@ Acknowledgments
This work is generously supported by the ONR grant 13983263 and the NSF grant CCF-2312205. The authors would like to thank Miao Lu and Wenhao Yang for pointing out an error in an earlier draft of this manuscript and their helpful comments. Z. R. is supported by Wharton Analytics. R. Z. is supported by RGC grant ECS-26210324.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Notation
[…]
Appendix B The ERM Solution
To solve the ERM problem, we follow Geman & Hwang (1982); Yadlowsky et al. (2022); Jin et al. (2022b) and adopt the method of sieves: consider an increasing sequence $\Theta_{1}\subset\Theta_{2}\subset\cdots$ of spaces of smooth functions, and let $(\widehat{\alpha}_{\pi}^{(k)},\widehat{\eta}_{\pi}^{(k)})=\underset{\theta\in\Theta_{n}}{\arg\!\min}~{}\mathbb{E}_{n}[\ell(X,Y(\pi(X));\theta)]$. In our case, we consider the following classes of sufficiently smooth functions. For $q_{1}=\lceil q\rceil-1$ and $q_{2}=q-q_{1}$ (where $q$ is the smoothness parameter), define the following function class for $\eta^{*}$:
[…]
where we denote the derivative $D^{d}=\sum_{l=1}^{p}\frac{\partial^{d_{l}}D}{\partial x_{l}}$. To ensure the non-negativeness of $\alpha^{*}_{\pi}$ in Proposition 2.5, we define the truncated function class $\Theta_{c}^{q}(\mathcal{X},\epsilon):=\{x\mapsto\max\{h(x),\epsilon\}:h\in\Theta_{c}^{q}(\mathcal{X})\}$ for the search of $\alpha_{\pi}$. Consequently, the function class we consider is $\Theta=\Theta_{c}^{q}(\mathcal{X})\times\Theta_{c}^{q}(\mathcal{X},\epsilon)$.
optimizers well, we need the true optimizer $(\alpha^{*}_{\pi},\eta^{*}_{\pi})$ to be sufficiently smooth in $x$. Convexity and stability are also desirable property of the loss function for learning the optimizers. In the next assumption, we present the regularity condition on the conditional distribution $\mathbb{P}_{Y\,|\,X}$ to ensure smoothness of optimizers.
Assumption B.1 (Smooth conditional reward distribution).
The conditional reward distribution $\mathbb{P}_{Y(a)\,|\,X=x}$ is smooth in $x$, i.e. for some $h\in\mathcal{X}$, $\mathbb{P}_{Y(a)\,|\,X=x+th}=\mathbb{P}_{Y(a)\,|\,X=x}+t\cdot\mathbb{P}_{h}$ for some measure $\mathbb{P}_{h}$ on $\mathcal{Y}$.
Taking practical terms into consideration, Assumption B.1 is reasonable as we assume the conditional distributions of $Y(a)$ are close for similar covariates, for any action $a\in\mathcal{A}$. Jin et al. (2022b, Appendix B.2) presents detailed discussion to justify smoothness of the optimizers under Assumption B.1.
Next, we would like to discuss the regularity conditions of the loss function $\ell$ in Equation (6) and its conditional expectation $\mathbb{E}[\ell(x,Y;\theta)\,|\,X=x,A=\pi(x)]$. In particular, we require stability of the loss function and its conditional expectation so that plugging in estimators of the optimizers will not cause large errors, which is a mild condition that can be satisfied under the first-order Taylor expansion condition. Readers can refer to (Van der Vaart, 2000; Jin et al., 2022b) for a more detailed discussion. Later, Lemma D.1 summarises the regularity conditions in formal terms and indicates that in our case, with KL-divergence and the loss function defined as in Equation (6), all the above regularity conditions are satisfied.
Appendix C Experiment Details
We let the context set $\mathcal{X}=\{x\in\mathbb{R}^{5}:\|x\|_{2}\leq 1\}$ to be the closed unit ball of $\mathbb{R}^{5}$ and let the action set to be $[3]$; the rewards $Y(a)$’s are mutually independent conditioned on $X$ with $Y(a)\mid X\sim N(\beta_{a}^{\top}X,\sigma_{a}^{2})$, for $a\in[3]$. We choose $\beta$’s and $\sigma$’s to be
@@ Appendix C Experiment Details
The training dataset $\mathcal{D}_{\text{train}}=\{(X_{i},\pi_{0}(X_{i}),Y_{i}(\pi_{0}(X_{i})))\}_{i=1}^{n}$ is generated with a given behavior policy $\pi_{0}$ (unknown to policy learning algorithms), which chooses actions conditioned on context $x$ according to the following rules:
@@
We also generate 100 testing datasets, each with sample size 10,000. Each testing dataset $\mathcal{D}_{\text{test}}$ consists of i.i.d. draws of data tuple $\{(X_{i},Y_{i}(1),Y_{i}(2),Y_{i}(3))\}_{i=1}^{n}$, and is generated similarly to the procedure described above.
@@
Appendix D Deferred Proofs of the Main Results
D.1 Proof of Lemma 2.3
@@ Appendix D Deferred Proofs of the Main Results
where the function $f_{\text{KL}}(x)=x\log x$ represents the KL divergence function. In (D.1), the first constraint reflects that $L$ is an likelihood ratio, and the second constraint corresponds to the KL divergence bound.
For notational simplicity, let $\mathbb{E}_{x}$ be the shorthand of $\mathbb{E}_{P_{Y\,|\,X}}[\cdot\,|\,X=x]$. By Theorem 8.6.1 of (Luenberger, 1997), the Slater’s condition is satisfied and strong duality holds:
@@
D.2 Proof of Theorem 3.5
For notational simplicity, we drop the dependence on $P$ in $\mathbb{E}_{P}$ when the context is clear. The proof of Theorem 3.5 makes use of the following lemma, which establishes some useful properties of the optimizer $\bm{\theta}^{*}_{\pi}$. The proof of Lemma D.1 can be found in Appendix F.1.
Lemma D.1.
@@
By the first-order condition established in part (1) of Lemma D.1, we have
@@
where the last step is due to Equation (13). By Assumption 3.4, $\|\widehat{\bm{\theta}}^{(k)}_{\pi}-\bm{\theta}^{*}_{\pi}\|_{L_{\infty}}=o_{P}(1)$. Therefore, for any $\beta\in(0,1)$, there exists $N\in\mathbb{N}_{+}$ such that for $n\geq N$, $\|\widehat{\theta}_{\pi}^{(k)}-\theta^{*}_{\pi}\|_{L_{\infty}}\leq\min(\xi,\xi_{1})$. On the event that $\|\widehat{\theta}_{\pi}^{(k)}(x)-\theta^{*}_{\pi}(x)\|_{L_{\infty}}\leq\min(\xi,\xi_{1})$ by part (2) of Lemma D.1 and Jensen’s inequality, we have
@@
where the last step is due to part (3) of Lemma D.1. Combining the above results, we have that
@@
Above, step (i) follows from Jensen’s inequality and step (ii) from part (3) of Lemma D.1. Combining everything, we have that term (IV) is of rate $o_{P}(n^{-1/2})$.
@@
For any $a\in[M]$, if we take $\pi(x)\equiv a$, then by (1) of Lemma D.1,
@@
By (2) of Lemma D.1, for any $a\in[M]$, there exists a constant $\xi_{a}>0$ such that for any $\|\theta-\bm{\theta}^{*}_{a}(x)\|_{2}\leq\xi_{a}$
@@
By (3) of Lemma D.1, for any $a\in[M]$, there exists a constant $\xi_{1,a}>0$ such that for any $\|\bm{\theta}-\bm{\theta}^{*}_{a}\|_{L_{\infty}}\leq\xi_{1,a}$,
@@ D.3.2 The effective term
Above, the first inequality is because of the definition of $\sup$ and the second is due to the triangle inequality; the last step is due to the boundedness of $h$. Taking the supremum over all $\pi\in\Pi$ in (D.3.2), we have that
@@
By the bounded difference inequality (Wainwright, 2019, Corollary 2.21), for any $t>0$,
@@
D.3.3 Bounding the negligible term
@@
Applying Lemma D.2 with $c_{i}=c_{1}(X_{i})$, we have that
@@
where the last step follows from part (1) of Lemma D.1. By Assumption 3.4, for any $\beta\in(0,1)$, there exists $N_{3}\in\mathbb{N}_{+}$ such that when $n\geq N_{3}$,
@@
Then by Lemma D.2, we have
@@
D.4 Proof of Theorem 4.6
We first state somes results from (Si et al., 2023) that will be used in the proof. For any $p,q\in[0,1]$, define
@@
Lemma D.3 (Adapted from Lemma A17 of Si et al. (2023)).
@@
Note that our definition of $g_{\delta}(q)$ is slightly different from that in (Si et al., 2023), so we include the proof of Lemma D.3 in Appendix F.4 for completeness.
@@
where step (i) uses that $g$ is non-decreasing (c.f. Cauchois et al. (2024, Proposition 1)); in step (ii), $\xi\in((1-\Delta)/2,(1+\Delta)/2)$, and step (iii) follows from Lemma D.3.
@@
(D.4)$\geq\sqrt{\frac{d}{n}}\frac{d2^{d-1}}{15d2^{d+2}}=\frac{1}{120}\times\sqrt{\frac{d}{n\varepsilon}}.$
Appendix E Generalization to Identifiable Covariate Distribution Shift
We now extend our methodology to handle situations where shift in $X$ and $Y\,|\,X$ distributions are both present. We note that, in most practical cases, the decision maker has access to the covariates in the target environment, making the shift covariate distribution identifiable and estimable — it is therefore unnecessary to guard against the worst-case shift.
Method.
Formally, suppose we have access to a training dataset $\mathcal{D}$ collected in an environment $P$, and aims at learning a policy that behaves well in the environment $Q$. Here, we assume that $Q_{Y\,|\,X}\in\mathcal{P}(P_{Y\,|\,X},\delta)$ for some given radius $\delta>0$ but do not impose any constraints on $X$ distribution shift except that $Q_{X}$ is absolutely continuous with respect to $P_{X}$ and that the likelihood ratio is bounded. Letting $r(x)=\frac{dQ_{X}}{dP_{X}}(x)$, we can estimate $r$ with $X$ from $P$ and $Q$ using standard tools (by directly estimating the density ratio or by means of classification algorithms). For any policy $\pi$, the distributional robust policy value can be written as
$\mathcal{V}_{\delta}(\pi):=\mathbb{E}_{Q_{X}}\Big{[}\inf_{Q_{Y\,|\,X}\in\mathcal{P}(P_{Y\,|\,X},\delta)}\mathbb{E}_{Q_{Y\,|\,X}}\big{[}Y(\pi(X))\,|\,X\big{]}\Big{]}.$ (25)
Note that the inner expectation of (25) is the same as in (1). By Lemma 2.3, there is
[…]
$=\mathbb{E}_{Q_{X}\times P_{Y\,|\,X}}\Bigg{[}\frac{\mathbbm{1}\{A=\pi(X)\}}{\pi_{0}(A\,|\,X)}\bigg{(}\bm{\alpha}^{*}_{\pi}(X)\exp\Big{(}-\frac{Y(A)+\bm{\eta}^{*}_{\pi}(X)}{\bm{\alpha}^{*}_{\pi}(X)}\Big{)}+\bm{\eta}^{*}_{\pi}(X)+\bm{\alpha}^{*}_{\pi}(X)\delta\bigg{)}\Bigg{]}$
$=\mathbb{E}_{P}\Bigg{[}r(X)\frac{\mathbbm{1}\{A=\pi(X)\}}{\pi_{0}(A\,|\,X)}\bigg{(}\bm{\alpha}^{*}_{\pi}(X)\exp\Big{(}-\frac{Y(A)+\bm{\eta}^{*}_{\pi}(X)}{\bm{\alpha}^{*}_{\pi}(X)}\Big{)}+\bm{\eta}^{*}_{\pi}(X)+\bm{\alpha}^{*}_{\pi}(X)\delta\bigg{)}\Bigg{]}.$
The above expression ensures that the robust policy value is identifiable with the data accessible to the decision maker. With this representation, subsequent policy evaluation and learning are similar to the pure concept drift case, and we provide the adaptation below.
In addition to $\mathcal{D}$, we let $\tilde{\mathcal{D}}=\{X_{i}\}_{i=1}^{m}$ denote the covariates from environment $Q$, i.e., $X_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}Q_{X}$. Assume that $\lim_{m,n\rightarrow\infty}m/n=\gamma$. As before, we adopt a $K$-fold cross-fitting scheme, where we split both $\mathcal{D}$ and $\tilde{\mathcal{D}}$ into $K$ non-overlapping equally-sized folds. For $k\in[K]$, we use $\mathcal{D}^{(k+1)}$ and $\tilde{\mathcal{D}}^{(k+1)}$ to obtain $\widehat{\pi}_{0}^{(k)}$, $\widehat{r}^{(k)}$, and $(\widehat{\bm{\alpha}}^{(k)}_{\pi},\widehat{\bm{\eta}}_{\pi}^{(k)})$ as estimates of $\pi_{0}$, $r$, and $(\bm{\alpha}_{\pi}^{*},\bm{\eta}_{\pi}^{*})$, respectively; we then use $\mathcal{D}^{(k+2)}$ and $\tilde{\mathcal{D}}^{(k+2)}$ to obtain $\widehat{g}_{\pi}^{(k)}$ as an estimate for $\bar{g}_{\pi}^{(k)}$. The estimator of the $k$-th fold is
$\widehat{\mathcal{V}}^{(k)}_{\delta}(\pi)$ $=\frac{1}{\mathcal{D}^{(k)}}\sum_{i\in\mathcal{D}^{(k)}}\frac{\widehat{r}^{(k)}(X_{i})\mathbbm{1}\{A_{i}=\pi(X_{i})\}}{\widehat{\pi}_{0}^{(k)}(A_{i}\,|\,X_{i})}\cdot\big{(}\widehat{G}^{(k)}_{\pi}(X_{i},Y_{i})-\widehat{g}_{\pi}^{(k)}(X_{i})\big{)}+\frac{1}{|\tilde{\mathcal{D}}^{(k)}|}\sum_{i\in\tilde{\mathcal{D}}^{(k)}}\widehat{g}_{\pi}^{(k)}(X_{i}),$
and the final robust policy value estimator is $\widehat{\mathcal{V}}_{\delta}(\pi)=-\frac{1}{K}\sum_{k=1}^{K}\widehat{\mathcal{V}}^{(k)}_{\delta}(\pi)$. We then obtain the learned policy via
$\widehat{\pi}_{\text{LN}}=\underset{\pi\in\Pi}{\arg\!\max}~{}\widehat{\mathcal{V}}(\pi),$
where we apply the same computational trick as in the pure concept shift case.
Theoretical Guarantees.
We now extend the theoretical guarantees to the general case. Since the proof is similar to the pure concept drift case, the proof sketch is provided. As a prerequisite, we modify Assumption 3.4 to be:
Assumption E.1.
For any policy $\pi$, assume that for each $k\in[K]$, the estimators $\widehat{\pi}_{0}^{(k)}$, $\widehat{r}^{(k)}$, $\widehat{g}_{\pi}^{(k)}$, and the empirical risk optimizer $\widehat{\bm{\theta}}_{\pi}^{(k)}$ satisfy
$\|\widehat{r}^{(k)}/\widehat{\pi}_{0}^{(k)}-r/\pi_{0}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=o_{P}(n^{-\gamma_{1}}),~{}\|\widehat{g}_{\pi}^{(k)}-\bar{g}_{\pi}^{(k)}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=o_{P}(n^{-\gamma_{2}}),$
$\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}^{*}_{\pi}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=o_{P}(n^{-\frac{1}{4}}),~{}\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}^{*}_{\pi}\|_{L_{\infty}}=o_{P}(1),$
Theorems E.2 and E.3 establish the asymptotic normality of the policy value estimator and the regret upper bound.
Theorem E.2.
Suppose Assumptions 2.12.43.3, and E.1 hold. Additionally assume that $dQ_{X}/dP_{X}\leq C$, a.s., for some constants $C>0$, and that $\lim_{m,n\rightarrow\infty}m/n=\gamma$. For any policy $\pi:\mathcal{X}\mapsto[M]$, we have $\sqrt{n}\cdot\big{(}\widehat{\mathcal{V}}_{\delta}(\pi)-\mathcal{V}_{\delta}(\pi)\big{)}\stackrel{{\scriptstyle\textnormal{d}}}{{\rightarrow}}N(0,\sigma^{2}_{\pi})$, where
$\sigma^{2}_{\pi}=\textnormal{Var}\bigg{(}\frac{r(X)\mathbbm{1}\{A=\pi(X)\}}{\pi_{0}(A\,|\,X)}\cdot\big{(}G_{\pi}(X,Y)-g_{\pi}(X)\big{)}\bigg{)}+\gamma\cdot\textnormal{Var}(g_{\pi}(X)).$
Theorem E.3.
Suppose Assumptions 2.12.43.3E.1 hold. Additionally assume that $dQ_{X}/dP_{X}\leq C$, a.s., for some constants $C>0$, and that $\lim_{m,n\rightarrow\infty}m/n=\gamma$. For any $\beta\in(0,1)$, there exists $N\in\mathbb{N}_{+}$ such that when $n\geq N$, we have with probability at least $1-\beta$ that
$\mathcal{R}_{\delta}(\widehat{\pi}_{\textnormal{LN}})\leq\frac{C(\kappa(\Pi)+\sqrt{\log(1/\beta)}),}{\sqrt{n}}$
where $C>0$ is a constant independent of $n$ and $\Pi$.
Proof Sketch.
Recall that
$G_{\pi}(x,y)=\bm{\alpha}^{*}_{\pi}(x)\exp\Big{(}-\frac{y+\bm{\eta}^{*}_{\pi}(x)}{\bm{\alpha}_{\pi}^{*}(x)}\Big{)}+\bm{\eta}^{*}_{\pi}(x)+\bm{\alpha}_{\pi}^{*}(x)\delta~{}\text{ and }~{}g_{\pi}(x)=\mathbb{E}\big{[}G_{\pi}(X,Y(\pi(X)))\,|\,X=x\big{]}$
It can be checked that
[…]
$=\mathbb{E}_{P}\Big{[}\frac{r(X)\mathbbm{1}\{A=\pi(X)\}}{\pi_{0}(A\,|\,X)}\cdot\big{(}G_{\pi}(X,Y)-g_{\pi}(X)\big{)}\Big{]}+\mathbb{E}_{Q_{X}}\big{[}g_{\pi}(X)\big{]}.$
So if we define $\bar{\mathcal{V}}_{\delta}(\pi)=-\frac{1}{K}\sum^{K}_{k=1}\bar{\mathcal{V}}^{(k)}_{\delta}(\pi)$, with
$\bar{{\mathcal{V}}}^{(k)}_{\delta}(\pi)$ $=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\frac{r(X_{i})\mathbbm{1}\{A_{i}=\pi(X_{i})\}}{\pi_{0}(A_{i}\,|\,X_{i})}\cdot\big{(}G_{\pi}(X_{i},Y_{i})-g_{\pi}(X_{i})\big{)}+\frac{1}{|\tilde{\mathcal{D}}^{(k)}|}\sum_{i\in\tilde{\mathcal{D}}^{(k)}}g_{\pi}(X_{i}),$
we have by the central limit theorem that
$\sqrt{n}\Big{(}\bar{\mathcal{V}}_{\delta}(\pi)-\mathcal{V}_{\delta}(\pi)\Big{)}\stackrel{{\scriptstyle\textnormal{d}}}{{\rightarrow}}N(0,\sigma^{2}_{\pi}).$
As in the proof of Theorem 3.5, we can decompose the difference between $\bar{\mathcal{V}}_{\delta}^{(k)}(\pi)$ and $\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)$ as follows:
$\widehat{\mathcal{V}}^{(k)}_{\delta}(\pi)-\bar{\mathcal{V}}^{(k)}_{\delta}(\pi)=$ $\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\frac{r(X_{i})\mathbbm{1}\{A_{i}=\pi(X_{i})\}}{\pi_{0}(X_{i})}\big{(}\widehat{G}^{(k)}_{\pi}(X_{i},Y_{i})-G_{\pi}(X_{i},Y_{i})\big{)}$
$+\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\mathbbm{1}\{A_{i}=\pi(X_{i})\}\Big{(}\frac{\widehat{r}(X_{i})}{\widehat{\pi}_{0}(X_{i})}-\frac{r(X_{i})}{\pi_{0}(X_{i})}\Big{)}\big{(}\widehat{G}^{(k)}_{\pi}(X_{i},Y_{i})-\bar{g}^{(k)}_{\pi}(X_{i})\big{)}$
$+\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\mathbbm{1}\{A_{i}=\pi(X_{i})\}\Big{(}\frac{\widehat{r}(X_{i})}{\widehat{\pi}_{0}(X_{i})}-\frac{r(X_{i})}{\pi_{0}(X_{i})}\Big{)}\big{(}\bar{g}^{(k)}_{\pi}(X_{i})-\widehat{g}_{\pi}(X_{i})\big{)}$
$-\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\frac{r(X_{i})\mathbbm{1}\{A_{i}=\pi(X_{i})\}}{\pi_{0}(X_{i})}(\widehat{g}^{(k)}_{\pi}(X_{i})-g_{\pi}(X_{i}))+\frac{1}{|\tilde{\mathcal{D}}^{(k)}|}\sum_{i\in\tilde{\mathcal{D}}^{(k)}}(\widehat{g}^{(k)}_{\pi}(X_{i})-g_{\pi}(X_{i})).$ (26)
Following almost the same steps in the proof of Theorem 3.5, we can show that
[…]
$\qquad+O_{P}\bigg{(}\Big{\|}\frac{\widehat{r}^{(k)}}{\widehat{\pi}_{0}^{(k)}}-\frac{r}{\pi}\Big{\|}_{L_{2}(P_{X\,|\,A=\pi(X)})}\cdot\|\widehat{g}^{(k)}_{\pi}-g_{\pi}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}\bigg{)}+O_{P}\big{(}\|\widehat{g}^{(k)}_{\pi}-g_{\pi}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}\cdot n^{-1/2}\big{)}$
$=$ $o_{P}(n^{-1/2}).$
Taking a union bound over $k\in[K]$, we can conclude $\widehat{\mathcal{V}}_{\delta}(\pi)=\bar{\mathcal{V}}_{\delta}(\pi)+o_{P}(n^{-1/2})$, and therefore $\widehat{\mathcal{V}}_{\delta}(\pi)$ is asymptotically normal. The proof of Theorem E.3 follows similarly.
Appendix F Proof of technical lemmas
F.1 Proof of Lemma D.1